### v1

# Clustering Algorithms and RAG Enhancing Semi-Supervised Text Classification with Large LLMs

Abstract
This paper introduces an innovative semi-supervised learning approach for text classification, addressing the challenge of abundant data but limited labeled examples. Our methodology integrates few-shot learning with retrieval-augmented generation (RAG) and conventional statistical clustering, enabling effective learning from a minimal number of labeled instances while generating high-quality labeled data. To the best of our knowledge, we are the first to incorporate RAG alongside clustering in text data generation. Our experiments on the Reuters and Web of Science datasets demonstrate state-of-the-art performance, with few-shot augmented data alone producing results nearly equivalent to those achieved with fully labeled datasets. Notably, accuracies of 95.41% and 82.43% were achieved for complex text document classification tasks, where the number of categories can exceed 100.
1 Introduction
Often, fine-tuning large language models for specific tasks demands a considerable amount of data. However, usually the real industry datasets we encounter are not that labeled accurately and precisely, or such well-labeled datasets are difficult to obtain, or costly for people to perform human labeling. Semi-supervised text classification (SSTC), a hot subbranch of text mining, is dedicated to improving classification performance by leveraging a small amount of labeled data and a large amount of unlabeled data. Unfortunately, traditional approaches[1][2] targeting the semi-supervised text classification (SSTC) task typically entail a complex model structure, which proves to be prohibitively challenging for people to train and utlize such models.
Meanwhile, Large Language Models (LLMs) have achieved remarkable progress in the domain of artificial intelligence recently. These models have drawn widespread attention because of their outstanding language understanding and generation capabilities, and have demonstrated great potential in numerous application fields. Large LLMs can not only function as generative models but essentially also as probabilistic models that possess the capacity to output choice probabilities and make decisions. This can be achieved, for example, by assigning choice IDs such as A and B and calculating the logit of the last output word before the EOS token. Recent research indicates that large language models are excellent zero-shot and few-shot learners by incorporating examples and instructions into the prompt.[3]. Technologies such as Chain of Thought (CoT)[4] and Tree of Thoughts[5] have subsequently further demonstrated significant enhancements in multiple text classification[6][7] benchmark tests by enabling LLMs to output reasoning before the answer, or by conducting further ensembles, i.e self-consistency[8] among multiple answers. Thus, employing LLMs as models for SSTC tasks can benefit from their generalization and texture understanding ability. Moreover, the popularity of LLMs, along with numerous developed training and deployment frameworks, makes them more attractive contrast to traditional SSTC models.
On the other hand, when we encounter datasets that encompass professional domain knowledge, relying merely on the capacity of LLMs is inadequate. Retrieval Augmented Generation (RAG) [9], a method first introduced in 2020, gained popularity after the emergence of LLM. RAG offers benefits such as associating answers with professional knowledge without undergoing training, thereby enhancing the answer quality and reducing hallucinations. This is achieved by first retrieving relevant information from external knowledge, mainly through embedding models, and subsequently combining the retrieved information into the input prompt of LLMs to generate answers. There are a number of works aiming at improving RAG pipeline from various aspects such as query rewrite, document rerank, and post-retrieval processing[10]. The RAG framework proved to be highly suitable for text classification tasks. In this regard, [11] utilized retrieval-enhanced LLMs to generate text classification datasetes under zero-shot condition. Moreover, [12] proposed a Retrieval-Augmented framework to alleviate poor generalization issues existed in text classification taskes.
However, for methods like RAG or CoT that do not entail fine-tuning, the outputs might not always comply with the instructions in the prompt, while fine-tuning enables adjustments of the behaviors of LLMs, such as adhering to specific writing styles[10]. At the same time, adding content to the prompt and generating reasoning prior to answering leads to increased computation costs. For instance, as of July 2024, for GPT-4-o, it is \$5 per 1M input and \$15 per 1M output, and the cost for text-embedding-3-large is \$0.13 .This can be depleted rapidly when we have a large volume of data and if we aim to apply CoT and RAG to every data point. In production scenarios where there is a massive amount of data and the answer is required as soon as possible, a smaller yet quicker specialized model that directly outputs the answer is much preferred. While [13] reveals that LLMs can improve themselves through the utilization of CoT prompting and self-consistency to generate high-quality answers from unlabeled data and fine-tuning based on these generated answers. Studies such as [14] have further demonstrated that fine-tuning a model with fewer parameters through knowledge distillation from the CoT outputs of larger LLMs results in improved task performance on datasets like GSM8K. Still, techniques such as RAG and CoT are excellent options when we encounter circumstances where there is no labeled data or only a very limited amount of labeled data, or when we aim to generate high-quality data.
Given the recent progress in methods such as RAG and CoT, which have enhanced the outputs of LLMs, yet considering the limitations of these methods that make them more suitable for improving a fine-tuned model rather than being directly used in production, especially for text classification tasks. Also, given the undeniable fact that the real-world data we encounter is often duplicated, mislabeled, or completely lacking labels and fine-tuning a model for specific tasks requires considerable amount of well labeled data. Since human experts are costly, it is essential to select the most representative landmark data that can accurately reflect the characteristics of the entire dataset. [15] employed a lightweight deep learning algorithm to select the most informative samples for both image and text classification datasets. Where [16] proposed a dynamic data selection approach for selecting ’long tail’ words in different epochs of training. While clustering can also be employed as a tool for data selection, more studies have focused on using the clustering technique as a feature selection tool [17][18].
This study presents an innovative approach for fine-tuning LLMs in text classification tasks by integrating multiple distinct components within a pipeline. Firstly, it involves data selection through clustering algorithms. Subsequently, human experts carry out manual labeling, and these labels act as retrieval examples via the RAG and CoT pipeline from a larger-sized LLM to generate additional training samples. Thereafter, these training samples are utilized to fine-tune a smaller-sized LLM and iteratively focus on the segments where the model might have made incorrect decisions, with these segments being determined again by clustering. Our innovation can be summarized as follows:
Clustering
Through controlled experiments, we demonstrate that our technique, which employs clustering to select representative landmark data from unlabeled datasets and pick potentially wrongly judged data points, enhances the accuracy of the fine-tuned model when the same amount of labeled data is utilized compared with random selection and the method proposed by [15] in datasets such as 20 newsgroups. Moreover, to the best of our knowledge, we are the pioneers in using clustering techniques for data selection in text classification tasks. In this manner, the labeled data ensures diversity and can cover the majority of cases, thereby well reflecting the characteristics of the entire dataset with a small quantity of data. Compared with the traditional full-scale labeling method, this approach significantly reduces the workload and time cost of manual labeling while ensuring the data quality.
Agumenting
Our study presented a novel approach that utilizes RAG to select samples from high-quality labeled data and combines them in a CoT chain to generate high-quality augmented data. We hold that the improvement stems from RAG, allowing the model to learn data patterns and regularities from a few labeled landmark data to boost accuracy, and CoT, which inspires LLMs themselves to consider potential corner cases. Such generated data can further expand the dataset, enhancing its diversity and richness. Despite numerous studies focusing on RAG and CoT, to our best knowledge, no research has employed them for data augmentation. In our experiment, this method significantly enhances the classification accuracy.
Integration
Although numerous studies have concentrated on each minor aspect of text classification, few have elaborated in detail on the different stages of text classification tasks, such as data selection, preprocessing, labeling, data augmentation, and iterative training with updated augmented data. Due to the increased diversity and representativeness of these selected labeled data, the trained model can deal with various practical scenarios more proficiently, improving the generalization ability and accuracy of the model.
Simple
Our intuitive approaches are straightforward to implement and understand, and each step is based on open source packages. This compatibility makes our method more pragmatic than other methods that pose challenges for people in terms of checking and implementation. We also offer our code on GitHub. Through fine-tuning LLM models with relatively smaller parameters (e.g., 4B), we found that in contrast to the chain of thought approach with a significantly larger parameter size (e.g., 72B), the model’s response accuracy is higher and it requires fewer resources.
2 Literature Review
2.1 Text Data Preprocessing, Selection and Augmentation
In tasks involving document text processing, a standard preliminary step involves tokenization, where the document is broken down into a sequence of individual words. Following this, a transformation process is commonly applied to convert the textual data into numerical representations to enable further analysis such as clustering and classification. Techniques like Term Frequency-Inverse Document Frequency (TF-IDF)[19] and fixed-length embeddings[20] are predominant in this conversion. Embedding methodologies transform document words into a series of vector representations, which can be seamlessly integrated into LLMs for downstrain tasks. This direct translation facilitates a deep understanding of semantic relationships within the text. Conversely, frequency-based method like TF-IDF offer a complementary perspective from a statistical summary viewpoint within and across documents, and usually operates much faster for large text dataset. Also, since the advent of LLMs, individuals have begun to develop various methods that are capable of filtering high-quality data in accordance with their requirements. For instance, Nvidia released a 340B parameter reward model [21] to filter data, which calculates a five-dimensional score including Helpfulness, Correctness, Coherence, Complexity, and Verbosity to filter data. There are also research concerns[22] regarding sample selection bias, such as repeating redundant samples, picking outliers, or disproportionately selecting from a subset of classes and developing a lopsided representation.
Besides data selection, data augmentation is also an important aspect as the need for data is huge. Traditional data augmentation techniques typically involve operations like randomly swapping, inserting, deleting words, or make use of synonym context replacement, which are accomplished via word lists or leveraging a trained language model[23]. Additionally, customized augmentation such as simulating misspelled words and errors stemming from image-to-text conversions are available to address specific scenarios. [24] summarized different text augmentation techniques, including rule-based augmentation, graph-structured augmentation, style augmentation, adversarial augmentation, as well many others techniques for augment text data. For more task-specific augmentation that necessitates domain expertise or involves modifying documents that have specific styles, particularly when dealing with semi-structured texts where certain words cannot be replaced, a more nuanced approach is required. [25] explored various techniques for augmenting legal document data. [26] adopt a random word replacement with wordnet to augment email datasets. Since the prevalence of LLMs, some individuals have utilized ChatGPT to generate and augment data. For example, [27] adopts a two-stage augmentation strategy to augment clinical annotation data, by first apply ChatGPT to generate samples that maintain semantic coherence with its label, rephrasing each input sentence into six additional variants, all adhering to the original sentence’s label, then a BERT model trained on the original samples is used to validate the generation based on the classification results of the generated samples and filter the generated samples. However, their method were only applied on short sentence texts.
2.2 Clustering
From the standard K-means clustering to those enabling the automatic adjustment of cluster center numbers, clustering has long been a traditional unsupervised learning method. LMSC[28] utilized the complementary information from multiple views of the data. It reconstructed original data by multiplying different projection vectors and subsequently performed clustering on these subspaces. When the number of available documents is large, it is usually time-consuming to conduct all clustering simultaneously. LBPSO[29] adopts a neighborhood selection strategy, learning from individual historical best solutions and neighborhood best solutions, avoiding calculating a global solution at once and preventing the algorithm from getting entraped in local optima. [30] mentioned a local fuzzy one pass clustering method where the one pass stands for accomplishing the clustering process in a single iteration over the dataset, thus can handle streaming incremental data.
2.3 Semi-Supervised Text Classification(SSTC)
Latent Dirichlet Allocation (LDA) topic models are parameter-optimized probabilistic models that can represent the text content of unlabeled instances, making them suitable for SSTC tasks. [31] utilized a self-training method starting from the initial small labeled dataset and for each new incoming document, demonstrating better performance compared to traditional TF-IDF representation methods and supervised classifiers. On the other hand, [32] proposed a method concentrating on the set of non-examples that do not belong to any of the interested categories. This can be employed to approximate the prior distribution of the classification function, thereby enhancing the accuracy of methods such as AdaBoost in SSTC tasks. In addition, [33] proposed an iterative self-pretraining method based on two classifiers. One classifier labels a random set of unlabeled documents, while the other is further trained on both labeled and pseudo-labeled documents in each iteration. Confidence thresholds were introduced to filter pseudo-labels. The final classification of unlabeled data is accomplished through an ensemble prediction of the two classifiers. Experiments indicate that a higher F-1 score was achieved compared to six other methods in serverl text classification tasks.
2.4 Chain of thought and RAG
The Chain of Thought (CoT) method, initially introduced by [4], entails incorporating a sequence of intermediary reasoning steps that precede the formulation of conclusions. CoT promping has demonstrated a significant boost in accuracy across a wide range of different tasks. [34] provided a theoretical viewpoint about why prompting works. They replaced with the wrong label for the few-shot examples and find it barely hurt performance. Instead they found the distribution of the input text in the few-shot examples, the overall format of the few-shot examples, and the label space of the few-shot examples are the key factors contributing to performance. When employing CoT methodologies, the extraction of the final answer is typically done by parsing the generated text through regular expressions, which prone to extraction errors.[35] developed a 500 million parameter LLM specifically designed to extract the answer from the generated responses, and showed an improved accuracy over regular expressions. Further exploration by [36] found CoT provides 10% to 20% accuracy gains over standard prompting in most tasks, except tasks requiring external world knowledge, where it’s accuracy lag behand standard prompting. Retrieval Augmented Generation (RAG) [9] is a method focused on improving generation performance by fetch text information from external knowledge bases and incorporate these information in the prompt to produce output. Retrieval Augmented Generation (RAG) [9] is a methodology designed to enhance the quality of text generation by integrating information retrieved from external knowledge sources into the generation process. This approach involves fetching relevant textual data into the input prompt, thereby producing more informed and contextually accurate outputs. Since the advent of RAG, studies[10] have exerted diverse efforts to enhance the RAG outputs. For example, incorporating evaluation modules to measure the appropriateness of employing RAG compared to direct text generation, or adopting distinct specialized workflows, such as exploiting knowledge graphs[37] or navigating through specific hierarchical levels of documents for retrieval, depending on the given context or scenario. Other examples include sophisticated paragraphs segmenting methods that reduce the irrelevant noise, transferring query to broaden or refine search intent, and the utilization of metadata filters to narrow down results. Additionally, hybrid retrieval models, which combine embeddings for semantic understanding with traditional tools like Elasticsearch and keyword matching, offer complementary strengths when employed in ensemble, or facilitate the switching of search strategies when dealing with large datasets to conserve time. Re-ranking or using different LLMs to refine selections, fine-tuning the retriever or generator, and iterative retrieval processes, where successive queries refine the result set based on the initial feedback, further enhance the accuracy and relevance of information retrieval. Collectively, these strategies improve search performance by adapting to user needs.
2.5 Finetuning
Despite the abundance of technical reports on various models that open-source their model weights, there are relatively less publicly available reports as well datasets on the techniques for training and fine-tuning language models. The training of LLMs typically involves stages such as pre-training, instruction fine-tuning, and preference fine-tuning[38][39]. Our focus lies on the instruction fine-tuning aspects. [40] investigated scaling the number of tasks, the number of data points, and the number of parameters. They found that multi-task instruction fine-tuning enhanced prediction accuracy: as the number of fine-tuned tasks increased, the prediction accuracy of an 8B parameter LLMs on the renowned Massive Multi-task Language Understanding (MMLU) dataset rose from 24.3% to 26.3%, 37.7%, 47.7%, and 49.3% for 9, 89, 282, and 1836 fine-tuned tasks respectively. This margin diminishes as the size of the model parameter becomes larger. [41] discussed about the dataset size with respect to the instruction fine-tuning performance. They selected instruction data of different scales from 200,000, 600,000, 1 million, to 2 million for comparison, and observed a steady overall improvement across different tasks as the trainable data increased. However, their training data was generated from ChatGPT with less than 200 high-quality human-labeled data for each task.[42] demonstrated that LLMs can effectively learn interaction styles or formats through fine-tuning with fewer than 1,000 carefully annotated training prompts and responses, without the need for data augmentation or other tricks. [43] discussed about parameter efficient fine-tuning approaches and compared the accuracy among different tunning methods such as vanilla fine-tuning, prompt-tuning, prefix-tuning, LoRA and adapter for over 100 tasks. They found that vanilla fine-tuning have the highest average results while prefix tuning, LoRA and adapter are not far behind. They also state that they did not observe an increase in performance with the increase of the efficient tuning parameters for at least small-sized LLMs. [44] compared parameter efficient finetunning of T0 model with GPT3 few shot prompt exampling and found a 6% higher accuracy is achieved. There also exist model structures that do not rely on the transformer structure but adopt the RNN structure, such as TTT and Manba[45][46]. Such methods might potentially perform better in scenarios where the document is long.
3 Proposed Method
Our methodology unfolds in three distinct phases:
3.1 Clustering Phase
In the initial phase, the emphasis lies on judiciously selecting a limited number of instances from a texture dataset for manual annotation. The dataset, denoted as $X=\{x_{1},x_{2},\ldots,x_{n}\}$, comprises individual documents where each $x_{i}$ symbolizes a sample from the original documents. The objective of this phase is to identify the most effective sampling strategy, $\pi$, under the practical constraint that only a subset $X_{\pi}^{m}$ of $m$ samples can be annotated due to the inherent limitations of human annotation resources. This subset $X_{\pi}^{m}$ of $m$, drawn from the dataset $X$ using the selection method $\pi$, is subsequently assessed by an evaluation metric $Q$, which we elect to be the several metrics that measure the performance of clustering using $X_{\pi}^{m}$. Desbribed in a mathematical way[47], the quest revolves around finding the best selection approach, $\pi^{*}$, which optimizes the chosen performance:
$\pi^{*}=\arg\max_{\pi}Q\left(X_{\pi}^{m},\pi\right)$
Several strategies exist for document data selection, primarily aimed at extracting domain-specific data from more general datasets [48][49]. These methods typically focus on selecting documents in a way that maintains a distribution or style similar to that of an existing target dataset, such as Wikipedia. To attain our specific objective, we utilized unsupervised clustering approaches to identify a greater quantity of smaller and more homogeneous clusters of documents, given the lack of a predefined target distribution. This method is more favorable for identifying representative landmark data that reflect the diversity of the entire dataset, rather than forming a few large clusters consistsing broader categories with a large number of documents. By forming these smaller, more concentrated clusters, we can utilize techniques such as exhaustively analyzing intra-cluster similarity or simply leverage LLMs as the example in Table 3.1 shown to infer and guide the selection of a single, high-quality, representative, most exemplary sample within each cluster. We can then denote the representative landmarks as:
$(X^{m},Y^{m})$
Where $Y^{m}$ are assumed to be annotated by human experts after obtained $X^{m}$.
Table 3.1: Prompt to Choose Representative Landmarks for Each Cluster You have been provided with a set of similar documents, each indexed by a number. Your task is to identify the most representative example from this cluster of documents. Please carefully analyze the given documents and select one document that best captures the common essence and characteristics of the samples. The selection should emphasize the representativeness and relevance of the chosen sample to the category, so that it can serve as a reference for labeling the entire cluster.
Documents in the cluster:
1-. {Document 1}……
2-. {Document 2}……
3-. {Document 3}……
4-. {Document 4}……
$\vdots$
Please choose one document that could best serve as a reference for labeling the entire cluster, and return only the index number of your selection, in format such as [0], [1], etc.
Answer:
3.2 Data Augmentating Phase
During the data augmentation phase, we employ a variety of data augmentation strategies to generate additional training instances. These strategies include using WordNet for word replacement, leveraging LLMs to rewrite documents, employing a combination of CoT reasoning with RAG to predict labels for each unlabeled sample, and directly generating data using RAG. Given the original unlabeled dataset $X$ and the labeled subset $(X^{m},Y^{m})$, our objective during the augmentation phase is to expand the labeled dataset through a range of augmentation techniques denoted as $G$:
$D=\{(x,y):(x,y)\sim G(X,X_{\pi}^{m},Y^{m})\}$
where $x$ stands for an augmented sample, $y$ denotes its corresponding label, and $D$ represents the final augmented dataset used for further model fine-tuning. For augmentation using WordNet and LLMs rewrite methods, only $X_{\pi}^{m}$ is utilized for augmentation purposes, while the corresponding labels $Y^{m}$ remain unchanged. In the case of random replacement using WordNet, a process akin to that described in [26] is adopted, where the WordNet [52] lexical database is used to identify a fixed number of the closest synonyms, and these are then randomly replaced with each other. For LLMs rewrite augmentation, we adopt the prompt strategy as outlined in Table 3.2.
Table 3.2: Prompt For LLMs Document Rewrite *Task Description: Rewrite the following text in English, maintaing the original meaning but using different words and sentence structures. The new version should be clear and concise, and it should not alter the core message of the original text. *Original Text: {The document upon which augmentation will be based} Rewritten Text:
Table 3.3: Prompt For RAG CoT Document Labeling *Task Description: You are provided with a set of similar documents. Your task is to predict the label for the target document, using the labeled document examples as references. *List of Available Labels: {List of available labels, seperated by comma} *Reference Labeled Documents: {A set of labeled landmarks from the nearest five clusters} *Target Document for Prediction: {The document upon which augmentation will be based} *Task: Predict the label for the target document. Please provide your reasoning before asssigning the label. *Format: Thought: [Your thoughts] Label: [Your assigned label] Answer:
For RAG-CoT based labeling of unlabeled samples and generating data using RAG, all $X,X_{\pi}^{m},Y^{m}$ are utilized. Specifically, when employing RAG CoT labeling, we leverage the cluster memberships established during the initial phase to improve our data labeling strategy. For each of the smaller clusters identified, we use LLMs to predict labels for each sample both individually and collectively. Collective predictions allow the LLM to consider the context jointly with other similarly labeled samples from nearby clusters, while CoT judgments help to enhance the accuracy of these predictions. With the details provided in Table 3.3.
Table 3.4: Prompt For RAG Dataset Augmentation *Task Description: You are provided with a set of similar documents, some of which are labeled and others are not. Your task is to generate a sample document based on the primary document, using both the labeled and unlabeled documents as references. *List of Available Labels: {List of available labels, seperated by comma} *Reference Labeled Documents: {A set of labeled landmarks from the nearest five clusters} *Reference Unlabeled Documents: {A set of unlabeled documents within the same cluster} *Primary Document for Augmentation: {The document upon which augmentation will be based} *Task: Using the labeled and unlabeled documents as a guide, create a new document based on the primary document and assign it the appropriate labels from the available list. *Document Format: Content: Label: [Your assigned label] *Generated Example:
For RAG document generation, we implement a refined approach that leverages both labeled landmarks from similar neighborhood clusters and unlabeled samples within the same clusters, as demonstrated in Table 3.4. For each original document, we identify the top five most probable clusters to which it might belong and select labeled landmarks from these clusters. We then randomly choose three unlabeled samples from the same cluster, combine them with the five labeled landmarks from neighboring clusters, and include the original document itself to construct a diverse prompt. This process does not require incorporating all documents within the cluster, and the random selection of reference documents enhances the diversity of the generated content. As a result, the average and maximum lengths of the prompts were reduced.
3.3 Phase 3: Fine-Tuning using LLMs
In the concluding phase of our project, we leverage the fine-tuning capabilities of pre-trained LLMs to enhance their performance. Fine-tuning LLMs allows us to produce more consistent answers compared to CoT methods and enables smaller models by parameter count to achieve performance comparable to larger models. The prompts designed for training the LLMs are outlined in Table 3.5:
Table 3.5: Prompt For LLM Training Assign tags for the following {subject} Document: {Document} Answer: [label1, label2, … ]
For label prediction tasks, the prompt structure is presented in Table 3.6:
Table 3.6: Prompt For Label Prediction Assign tags for the following {subject} Document: {Document} Answer:
During the training phase, the model is instructed to output labels in the ‘[label1, label2, …]‘ format. An explicit stop condition is set to ensure that the model begins generating with the opening bracket ‘[‘ and ceases text generation immediately after the closing bracket ‘]‘. Furthermore, to guarantee that the model only produces desired responses, we tokenize the available labels and constrain the model to output only the token IDs associated with these labels. For single-labeled data, the format ‘[label]‘ can be utilized; for multi-labeled data, if the training data are organized by importance from highest to lowest, LLMs will learn to output labels in the sequence ‘[label1, label2, …]‘ reflecting this hierarchy. In hierarchical labeling scenarios, the training data can be structured as ‘[tier1, tier2, …]‘, where ‘tier1‘ labels come from a predefined list of tier1 labels, and ‘tier2‘ labels follow accordingly, thus making the model’s output highly adaptable.
These fine-tuning procedures are conducted within the Llamafactory framework [53]. Our primary goal is to identify the optimal combination of training datasets that yield the best performance metrics when evaluated against test data. We consider various training dataset configurations, such as using landmark data alone, augmenting landmark data with WordNet, LLMs rewriting techniques, or employing RAG. Additionally, we seek to determine the smallest model size that meets performance requirements while minimizing computational resources. The study includes a comparative analysis between fine-tuning results and those obtained by combining RAG and CoT approaches. Specifically, for the CoT and RAG methods, the Qwen2.5-72b model is utilized, whereas for the fine-tuning experiments, the Qwen2.5-0.5B model are employed. This experimental setup enables us to show how different types of training datasets impact performance on specific tasks, offering valuable insights for real-world applications. Notably, all training was conducted within the constraint of an 80 GB GPU memory limit across two A100 GPUs.
4 Metric
4.1 Clustering
4.1.1 Homogeneity
The Homogeneity[55] metric assesses whether each cluster contains data points exclusively from a single class, reflecting the purity of the clustering relative to a ground truth. Mathematically, homogeneity is defined as:
$H(L_{true},L_{pred})=1-\frac{H(L_{pred}\mid L_{true})}{H(L_{true})}$
where $H(L_{pred}\mid L_{true})$ is the conditional entropy and $H(L_{true})$ is the entropy of the true labels. This metric is invariant under label permutations and non-symmetric; swapping true and predicted labels yields the completeness score. A score of 1 signifies perfect homogeneity.
4.1.2 NMI
Mutual Information (MI) quantifies the amount of information one clustering provides about another, making it a useful measure of similarity between two clusterings of the same dataset. Given two clusterings $U$ and $V$, MI is defined as:
$I(U;V)=\sum_{u\in U,v\in V}\frac{|u\cap v|}{N}\log\left(\frac{N\cdot|u\cap v|}{|u|\cdot|v|}\right)$
To address the potential bias due to cluster size differences and random cluster alignments, Normalized Mutual Information (NMI) standardizes MI. NMI scales MI by the average entropy of the two clusterings, ensuring the measure ranges from 0 to 1:
$\text{NMI}(U;V)=\frac{I(U;V)}{\text{avg}(\text{H}(U),\text{H}(V))}$
Here, $\text{H}(X)$ is the Shannon entropy of clustering $X$. An NMI of 1 indicates perfect agreement, while values close to 0 suggest near-independence, making NMI a valuable tool for evaluating clustering algorithms.
4.1.3 Silhouette Score
The Silhouette Score, $s(i)$, measures how much an object $i$ is closer to its own cluster than to others. It is defined as:
$s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$
Here, $a(i)$ is the mean distance between $i$ and all other points in its cluster, and $b(i)$ is the smallest mean distance between $i$ and points in any other cluster. Scores range from -1 to 1, with higher values indicating better cluster matches and values near 0 suggesting ambiguity.
4.2 Augmentation
4.2.1 Jaccard Similarity
The Jaccard Similarity is a measure that indicates the closeness of two sets by comparing the size of their intersection to the size of their union. It is defined as:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
Where $|A\cap B|$ is the number of elements in the intersection of sets $A$ and $B$, and $|A\cup B|$ is the number of elements in their union. The Jaccard Similarity coefficient ranges from 0 to 1, where a score of 1 means the sets are identical, and 0 indicates no overlap.
5 datasets
To conduct experiments and compare our work with others, we utilized several publicly available text datasets: The WOS-46985[56] dataset which is a subset of the Web of Science Dataset that comprises approximately 2.8 million scientific papers, the subset contains the abstract of 46,985 published paper, seven broad categories, including computer science, electrical engineering, psychology, mechanical engineering, civil engineering, medical sciences, and biochemistry, and 134 specific subfields. The charastric of published paper abstract made the length of the text longer than classification with just one sentence. The ApteMod version of the Reuters-21578[57] corpus includes 10,788 documents from Reuters financial newswire service dating back to 1987, grouped into 90 distinct topics such as corn, crude oil, earnings, gold, interest, and trade. Each document is labeled with one or more categories. The corpus is commonly used for training and testing machine learning models, particularly for multi-label text classification tasks.
The dataset is structured into three distinct subsets: 20% for testing, 30% for validation, and 50% for training. Within the 50% training subset, the true labels are hidden, but full access to the text context is provided for clustering, retrieval, and data augmentation. It is important to note that the 30% validation and 20% testing sets are strictly reserved for evaluating the performance of the models and not be used for data augmentation or clustering. This ensures an unbiased assessment of our procedure’s capabilities.
Table 1: Comparison of Text Datasets
Dataset WOS-46985 Reuters-21578
Dataset Size 46,985 10,788
Category Number 134 90
Average word count 199.8 164.9
Train Size 23,492 5,394
Validation Size 14,095 3,883
Test Size 9,398 2,158
6 Experiment
The initial phase of our experimental methodology involved applying clustering algorithms to the distinct Reuters and Web of Science (WOS) datasets. To validate these methods, we utilized the validation set, leveraging known labels to evaluate clustering performance, whereas we assumed no access to the labels of the training set. From this training set, a carefully selected proportion—300 for the Reuters dataset and 600 for the WoS dataset—was chosen to reveal the true labels. This selection process occurred during Phase 1, where specific data points were identified as landmarks—critical examples particularly valuable for labeling. Specifically, we compared embeddings and TF-IDF models across different cluster sizes and assessed the impact of various clustering methods using metrics such as Normalized Mutual Information (NMI), Homogeneity, and Silhouette Score. Subsequently, we extended our analysis to the training data, reinforcing the findings from the validation phase.
There are a variety of clustering techniques and feature representation methods in our disposal. For the data selection phase, we selected four clustering algorithms and two feature representation strategies that are particularly well-suited for scenarios with large numbers of samples and clusters. The chosen clustering methods include Bisecting K-Means, Balanced Iterative Reducing and Clustering Using Hierarchies (BIRCH)[50], Hierarchical Clustering and GaussianMixture Models. For feature representation, BGE-m3 Embedding[51] and TF-IDF techniques were chosen. The BGE-m3 method embeds up to 8192 tokens into 1024-dimensional vectors. For TF-IDF, we selected severak different max features terms based on frequency scores to address sparsity in the TF-IDF matrix. To evaluate the suitability of these combinations for our specific needs, we measure their performance using several evaluation metrics. Given that we possess the ground truth labels and our goal is to use clustering to identify more representative samples rather than just assigning predicted cluster labels, we are particularly focused on the Homogeneity score, which measures the extent to which each cluster contains only members of a single class. In addition to homogeneity, we also employ the Normalized Mutual Information (NMI) and Silhouette Score to assess the quality of the clustering.
Table 2: Comparison of Embedding Methods with TF-IDF Methods under Different Max-Features Settings for Reuters-21578 Dataset, using Hierarchical Clustering Methods, under the constrain of 300 clusters.
Tf-idf Max Features Homogeneity NMI Silhouette Score
256, TF-IDF 0.7878 0.4784 0.0726
512, TF-idf 0.7929 0.4832 0.0723
1024, TF-idf 0.8104 0.4936 0.0657
2048, TF-idf 0.8003 0.4916 0.0641
4096, TF-idf 0.7895 0.4875 0.0596
-, Embedding 0.8338 0.5123 0.0579
To evaluate the effectiveness of different feature extraction methods (TF-IDF and Embedding) on text clustering, we conducted experiments using the Reuters-21578 dataset with varying dimensions of TF-IDf max feature spaces (256, 512, 1024, 2048, 4096) and Hierarchical clustering algorithm, under the constrain of 300 clusters. Table 2 summarizes the results in terms of homogeneity, Normalized Mutual Information (NMI), and Silhouette Score. For TF-IDF based clustering on the Reuters-21578 dataset, the optimal homogeneity and NMI are achieved with 1024 max features. Deviating from this setting slightly reduces these metrics. The Silhouette Score consistently declines as the number of features decreases. Embeddings outperform TF-IDF with hierarchical clustering, achieving the highest Homogeneity (0.8338) and NMI (0.5123). Additionally, embedding methods do not require tuning for max features, unlike TF-IDF methods. Other clustering methods yield similar results.
Table 3: Comparison of Embedding Methods with TF-idf Methods under Different Max-Features Settings for WOS-46985 Dataset, using Hierarchical Clustering Methods, under the constrain of 600 clusters.
Tf-idf Max Features Homogeneity NMI Silhouette Score
512, TF-idf 0.5058 0.4432 0.0226
1024, TF-idf 0.5455 0.4779 0.0306
2048, TF-idf 0.5573 0.4907 0.0277
4096, TF-idf 0.5547 0.4925 0.0212
8192, TF-idf 0.5453 0.4879 0.0136
-, Embedding 0.6530 0.5677 0.0116
As shown in Table 3, embedding methods consistently outperform TF-idf methods across all different TF-IDF maximum feature settings for the WOS-46985 dataset. This suggests that embedding methods are better at capturing the deep semantic structures within the text, which is crucial for the classification criteria of the WOS-46985 dataset. Furthermore, the optimal vocabulary size for TF-idf was found to be between 2048 and 4096, indicating that scientific documents require a larger vocabulary for accurate representation compared to Reuters news articles. The lower Homogeneity and NMI scores for TF-idf also suggest a more challenging classification problem. Detailed grid search results confirmed that embedding methods achieve higher homogeneity and NMI scores, leading us to adopt embedding methods for our clustering approach.
Table 4: Comparison of Clustering Methods with Different Cluster Sizes on the Reuters Dataset
Clustering Method Number of Clusters Homogeneity NMI Silhouette Score Random 100 0.1195 0.0720 / Random 300 0.2679 0.1389 / Random 600 0.4042 0.1934 / Brich 100 0.7299 0.5194 0.0354 Brich 300 0.8147 0.5119 0.0446 Brich 600 0.8593 0.5021 0.0554 Bisecting K-Means 100 0.7101 0.4941 0.0230 Bisecting K-Means 300 0.7916 0.4777 0.0127 Bisecting K-Means 600 0.8473 0.4702 0.0034 Hierarchical 100 0.7467 0.5335 0.0421 Hierarchical 300 0.8338 0.5123 0.0579 Hierarchical 600 0.8839 0.4963 0.0835 Gaussian Mixture 100 0.7513 0.5275 0.0389 Gaussian Mixture 300 0.8127 0.4934 0.0408 Gaussian Mixture 600 0.8765 0.4944 0.0583
From Tables 4 and 5, it can be observed that Hierarchical Clustering and the Gaussian Mixture Model (GMM) outperform other clustering methods in most scenarios. While BIRCH and Bisecting K-Means are algorithms designed for computational efficiency, Hierarchical Clustering and GMM require more processing time, with GMM being the most computationally demanding. In practice, the output of BIRCH is often used as an initial point for hierarchical clustering or GMM. In addition to the encoding method employed, the size of the cluster numbers also impacts the homogeneity and normalized mutual information (NMI) scores. As the number of clusters increases, both homogeneity and NMI scores tend to rise. However, given that the objective of our clustering process is to identify landmark points for labeling, the number of clusters we can manage is constrained by the availability of human resources for labeling. For the Reuters dataset, we selected 300 clusters, whereas for the WOS dataset, we chose 600 clusters.
Table 5: Comparison of Clustering Methods with Different Cluster Sizes for the WOS Dataset
Clustering Method Number of Clusters Homogeneity NMI Silhouette Score Random 150 0.1565 0.1539 / Random 300 0.2551 0.2345 / Random 600 0.3713 0.3206 / Random 1200 0.4962 0.4044 / Birch 150 0.5141 0.5099 0.0067 Birch 300 0.5785 0.5369 0.0087 Birch 600 0.6420 0.5609 0.0097 Birch 1200 0.7087 0.5859 0.0138 Bisecting K-Means 150 0.4738 0.4685 0.0051 Bisecting K-Means 300 0.5338 0.4943 0.0002 Bisecting K-Means 600 0.5940 0.5168 -0.0077 Bisecting K-Means 1200 0.6642 0.5436 -0.0176 Hierarchical 150 0.5240 0.5202 0.0073 Hierarchical 300 0.5865 0.5440 0.0081 Hierarchical 600 0.6530 0.5677 0.0116 Hierarchical 1200 0.7225 0.5917 0.0190 Gaussian Mixture 150 0.5516 0.5452 0.0206 Gaussian Mixture 300 0.6017 0.5564 0.0177 Gaussian Mixture 600 0.6532 0.5681 0.0130 Gaussian Mixture 1200 0.7033 0.5774 0.0092
We also observed that the Gaussian Mixture Model (GMM) exhibits better performance compared to the Hierarchical Model when the number of clusters is smaller. However, as the size of the clusters increases, the Hierarchical Model surpasses GMM in turns of NMI, Homogeneity and Silhouette Score. An additional advantage of the Gaussian Mixture Model is its probabilistic nature, which enhances the interpretability of cluster membership. GMM provides probabilities for each data point belonging to different clusters, thereby allowing us to identify multiple potential cluster affiliations for each data point. This feature makes GMM particularly suitable for our subsequent step involving rag retrieval. It’s important to note that both Gaussian Mixture Models (GMM) and hierarchical clustering can generate a list of the most likely clusters for each sample. GMM provides a probabilistic framework, assigning a probability density to each sample across all clusters, whereas hierarchical clustering offers a nested structure of subclusters within larger clusters, which can be used to identify the nearest clusters for a given sample.
Figure 1: GMM Cluster Size Distribution for Reuters-21578 and WOS-46985 Datasets
We then proceed to analyze the training dataset, as illustrated in Figure 1. The resulting clusters exhibit a distribution that is somewhat less skewed than anticipated. Typically, clusters are unevenly distributed, with a majority of elements concentrated within a few large clusters, while many smaller clusters contain only one or two members. For the Reuters dataset, which comprises a total of 300 clusters, the largest cluster contains 100 elements, while the smallest includes just 3 members. The median size of the clusters is 14, and the 95th percentile is at 5 members. The silhouette score for this dataset is 0.0472. For the WOS dataset, consisting of 600 clusters, the largest cluster has 105 elements, and the smallest has only 2 members. The median size of the clusters is 37, and the 95th percentile is at 15 members. The silhouette score for the WOS dataset is 0.1523. Note that metrics requiring true labels are not available, as we assume we only have access to a few labeled training data points in a semi-supervised setting.
[…]
Following the selection and revelation of the true labels of the most representative landmarks from each cluster, we proceeded to utilize all available unlabeled data, along with the labeled landmarks ($X$, $X_{\pi}^{m}$, $Y^{m}$), to generate additional training data. For the RAG generation using prompt as shown in Table 3.4, the average prompt length was 1,836.25 words for the Reuters dataset and 2,260.34 words for the Web of Science (WoS) dataset, with the longest prompts comprising 5,607 and 3,573 words, respectively. For each document, we utilized the prompt to generate three unique samples by varying the randomly selected reference documents. Subsequently, we applied regular expressions to parse out the relevant context and labels from the generated texts. Out of the 16,182 documents generated for the Reuters dataset, 15,473 successfully underwent the regular expression-based extraction of context and labels. Similarly, from the 70,476 documents generated for the WoS dataset, 62,886 documents passed through the extraction process. Additionally, for the rewriting task using large LLMs, as detailed in Table 3.2, we generated ten samples for each of the landmark documents within the datasets. This was conducted with the hyperparameter temperature set to 0.3, yielding a total of 3,000 rewritten documents for the Reuters dataset and 6,000 for the WoS dataset. Since these generated samples did not require label extraction, they all successfully passed the regular expression filtering phase. It is noteworthy that the prompt lengths for the rewriting tasks were significantly shorter compared to those used for RAG generation, averaging 254.95 words for the Reuters dataset and 251.4 words for the WoS dataset.
Figure 2: Comparison of Category Distributions Before and After Augmentation with Three Samples per Document in the Training Set.
For the Reuters dataset, we configured the LLMs to produce multiple labels for each document, with individual labels delineated by commas. In contrast, for the WoS dataset, the LLM was instructed to generate a hierarchical labeling structure, beginning with the domain and then specifying the sub-domain, with each level separated by commas. Subsequently, the generated documents were evaluated against their original counterparts using a suite of performance metrics. While the labels for LLMs rewrite are kept the same as the original labels, we investigated the labeling of RAG-generated examples, as depicted in Figure 2. The Reuters dataset originally comprised 90 distinct labels, which were expanded to 443 unique labels through the use of LLMs. This expansion partly results from LLMs generating hallucinatory labels due to lack of fine-tuning for precise label generation. This process utilized 300 labeled landmarks, with 59 labels being covered. Despite this, 93.49% of documents in the augmented dataset still carry labels from the original 90, covering 85 of them. The training set includes 87 out of the original 90 labels. Compared to the original dataset, the augmented version shows increased category sizes. While only 21 categories had over 50 samples originally, now 52 categories exceed this threshold. It is expected that this number will increase further as more samples are augmented. However, there are 9 instances where the number of augmented documents is fewer than the original, often due to rare terms such as "jet," "cotton-oil," "dfl," "cpu," and "naphtha," which appeared only one or two times in the training data. Specifically, the "lei" category, representing the Leading Economic Index, saw no augmentation despite having 8 instances in the training set, likely because LLMs failed to recognize "lei" without context. On average, the augmentation process resulted in 268.14% more documents per category.
For the Web of Science (WoS) dataset which comprising 7 domains and 134 sub-areas, the augmented dataset introduces 11 domains, with 99.90% of the new domains aligned with the original 7. Regarding sub-areas, the augmented dataset generates 2,765 new sub-areas, of which 82.63% remain within the 134 sub-areas of the original WoS dataset, and 133 of the 134 original sub-areas are included. There are 7 cases where the augmented data has fewer samples compared to the original data. Upon examination, the least number of augmented documents, "medical, polycythemia vera" consists of 46, compared to 76 in the training set. The proportionate least is "psychology, attention," with 216 in the training set and only 52 in the augmented data. On average, the augmented data increased the content per category by 178.99%.
Figure 3: Comparison of Smoothed Word Lengths per Document Between Original and Augmented Documents in the Reuters and WoS Datasets
The average word count for the original context in the Reuters dataset is 128.5 for the training set, whereas the augmented data exhibits an average length of 167.4 words for RAG augmentation. For rewrite augmentation, the bases for augmentation are the landmarks, average length have 177.39 and 177 words for the augmentation. Regarding the WoS dataset, the initial documents within the training set have an average word count of 198.3, while the augmented documents extend to an average of 218.3 words. For rewrite augmentation, the bases for augmentation are the landmarks, average length have 208.40 and 191.24 words for the augmentation. It is observed that for RAG augmentation, when the original documents were relatively short, the synthesized versions tended to be more expansive. Conversely, if the original documents were already quite lengthy, the synthesized documents only showed a marginal increase in length on average. While for LLM rewrite, the generated documents are on average a little bit shorter than original. These observations are visually depicted in Figure 3. When considering the average character length per word, there is a slight increase of 1.4% for the Reuters dataset and 2.3% for the WoS dataset; these increments are statistically significant, as indicated by a very small standard error. A small indication that LLMs tends to use complex words.
Reuters WOS
Among Gen. Orig. vs Gen. Emb. Sim Among Gen. Orig. vs Gen. Emb. Sim
WordNet 94.57% 97.75% 97.75% 97.13% 98.52% 98.52%
LLM Rewrite 78.36% 47.79% 93.51% 76.32% 46.69% 96.35%
RAG Augment 49.12% 42.31% 92.17% 40.91% 36.08% 90.84%
Table 6: Comparison of different methods on Reuters and WOS datasets.
Figure 4: Impact of Text Augmentation on Jaccard Similarity (Vocabulary Overlap) among Generated Samples: The figure illustrates that from RAG augmentation to LLMs rewrite to traditional WordNet random synonym replacement, the vocabulary overlap decreases, indicating the generation of more diverse samples.
To evaluate the impact of various text augmentation techniques on the Reuters and WOS datasets, we conducted a comprehensive analysis focusing on the similarity between the original and augmented documents, as well as the diversity among the augmented samples derived from the same source document. Two primary metrics were used to quantify these aspects: Jaccard Similarity, which measures the degree of vocabulary overlap between sets of words, and Embedding Similarity, which assesses semantic similarity. A score closer to 1 indicates a higher degree of similarity. Our findings, summarized in Table 6 and illustrated in Figure 4, show that different augmentation methods yield varying levels of vocabulary overlap and semantic similarity:
  • WordNet Random Synonym Replacement: This method produced the highest vocabulary overlap, with Jaccard Similarity scores of 94.57% for the Reuters dataset and 97.13% for the WOS dataset, based on 10 samples generated from each landmark document, indicating that this technique preserves much of the original text’s structure and vocabulary. Longer documents generally exhibited higher similarity when augmented using WordNet
  • LLMs Rewrite: This method provided a balanced approach, achieving Jaccard Similarity scores of 78.36% for the Reuters dataset and 76.32% for the WOS dataset, based on 10 samples generated from each landmark document. It effectively introduced new variations while maintaining the core essence of the original content.
  • RAG Data Augmentation: This method resulted in the most diverse set of samples, with the lowest Jaccard Similarity scores of 49.12% for the Reuters dataset and 40.91% for the WOS dataset, calculated among the 3 samples generated using the landmark documents and other RAG components. The ability of this method to integrate diverse and randomly selected supporting materials led to significant modifications, thereby enhancing the diversity of the generated content.
Figure 4 further illustrates the impact of text augmentation on Jaccard Similarity (vocabulary overlap) among generated samples. From RAG augmentation to LLMs rewrite to traditional WordNet random synonym replacement, the vocabulary overlap decreases, indicating the generation of more diverse samples. When comparing the Jaccard Similarity between generated samples and the original documents, RAG augmentation achieved the lowest scores, averaging 42.31% for the Reuters dataset and 36.08% for the WOS dataset. Despite the lower vocabulary overlap, the embedding-based similarity remained high, with scores of 92.17% for the Reuters dataset and 90.84% for the WOS dataset. This suggests that while RAG augmentation introduces significant changes, it maintains the core meaning and semantic coherence of the original text. The RAG method stands out for its ability to introduce new information while preserving the original context, making it ideal for data augmentation in machine learning and content diversification in NLP tasks. The high embedding similarity scores further validate its effectiveness in maintaining semantic integrity despite significant textual modifications.
7 LLM Finetunning
7.1 Landmark vs Random
In our experiments, we utilized augmented data to train a Qwen2.5 0.5B language model to evaluate the performance of different augmented datasets. For the Reuters dataset, we evaluated the model using three metrics: Part Match, All Match, and In Right Order, based on a test set of 2,158 samples. These metrics are defined as follows:
  • Part Match: At least one of the generated tags aligns with one of the actual tags.
  • All Match: All generated tags perfectly match the true tags.
  • In Right Order: All tags match and are presented in the correct order of importance.
For the Web of Science (WoS) dataset, we evaluated the model using Domain Match and Area Match, which measure the accuracy of predictions at different hierarchical levels. The test set size was 9,398 samples. These metrics are defined as follows:
  • Domain Match: The generated domain matched with the true domain
  • Area Match: The generated area matched with the true area
To evaluate the effectiveness of our clustering techniques compared to random selection, we conducted a comparative analysis using two main types of data: landmark data and random data. The landmark data was carefully selected to represent meaningful and representative samples from the dataset, reflecting core themes. In contrast, the random data was chosen without specific criteria, serving as a baseline to gauge the added value of the clustering technique. To ensure a fair assessment of randomness, we performed three trials for the random data due to computational constraints. The datasets used in this study are detailed below:
  • Landmark Data: Selected landmarks believed to be indicative of the core themes within each dataset, 300 from Reuters and 600 from WoS. These samples were trained for 15 epochs, which yielded the highest performance.
  • Random Data: An equivalent number of samples (300 from Reuters and 600 from WoS) randomly chosen from the training set to serve as a control group. They were also trained for 15 epochs, achieving optimal performance.
  • Landmark-WordNet Data: The landmark data was augmented tenfold using WordNet, expanding the dataset to 3,000 samples from Reuters and 6,000 from WoS. Train for 1.5 epoch.
  • Random-WordNet Data: Similarly, the random data was also augmented tenfold using WordNet, resulting in 3,000 samples from Reuters and 6,000 from WoS. Train for 1.5 epoch.
Dataset Part Match All Match In Right Order
Trail 1, random data 88.92% 80.81% 74.28%
Trail 2, random data 88.13% 82.90% 76.08%
Trail 3, random data 87.85% 80.76% 74.23%
Landmark data 90.85% 82.35% 75.86%
Trail 1, random wordnet data 89.52% 82.25% 76.43%
Trail 2, random wordnet data 89.57% 83.45% 75.57%
Trail 3, random wordnet data 88.13% 81.74% 73.54%
Landmark wordnet data 93.05% 83.97% 78.82%
Table 7: Performance Comparison: Landmark vs. Random Data for Reuters.
The results, as summarized in Table 7 for the Reuters dataset, reveal several interesting insights. Initially, it might seem counterintuitive that in some trials, random data outperformed landmark data in terms of "All match" and "In right order" accuracy. However, this phenomenon can be attributed to the fact that when the data volume is small, more data about the majority classes is needed to improve model performance. During random selection, more samples are likely to fall into the majority class. Additionally, the test set contains more data representing common cases rather than edge cases. Therefore, in the initial stage of improving accuracy, more data describing the majority of the dataset is required to ensure the model generalizes well to typical scenarios. Despite this initial finding, the true value of the landmark data becomes evident when the datasets are augmented with WordNet. The simple augmentation process generated more samples for each category, which in turn significantly boosted the performance of the landmark data over the random data.
For instance, the Part Match metric for the Reuters landmark data improved from 90.85% to 93.05% after augmentation, representing a 2.2% enhancement in the model’s ability to generate at least one relevant tag. This is significantly higher than the 0.77% average improvement seen with random data. Similarly, the All Match and In Right Order metrics also showed substantial gains, increasing from 82.35% to 83.97% (1.62%) and from 75.86% to 78.82% (2.95%), respectively. In contrast, the average improvements for All Match and In Right Order with random data were only 0.99% and 0.31%. These results indicate that the clustering technique used to select the landmark data provides more meaningful and valuable insights for enhancing model performance. While random data can sometimes offer better initial performance, the strategic use of landmark data and semantic augmentation through resources like WordNet can lead to more robust and accurate tag generation.
Domain Match Area Match
Trail 1, random data 74.84% 47.13%
Trail 2, random data 77.25% 47.09%
Trail 3, random data 72.03% 46.10%
Landmark data 73.85% 47.48%
Trail 1, random wordnet data 76.99% 51.21%
Trail 2, random wordnet data 77.32% 50.57%
Trail 3, random wordnet data 77.30% 50.06%
Landmark wordnet data 77.35% 52.09%
Table 8: Performance Comparison: Landmark vs. Random Data for WOS.
Similarily, for the results shown in Table 8, the landmark-WordNet data achieved the highest accuracy in both Domain Match (77.35%) and Area Match (52.09%), with 3.5% and 4.61% increase. Without augmentation, the landmark data slightly underperformed compared to random data in both metrics. Similar to the reuter dataset, this could be because the landmark data selection focuses on capturing core themes within the dataset, which may not always align perfectly with the overall distribution of the majority of samples. In contrast, random data selection tends to better reflect the overall distribution of the dataset, especially when the sample size is small, making it more likely to cover the main categories. After applying WordNet augmentation, there is a significant performance improvement for both landmark and random data in both Domain Match and Area Match. Notably, the landmark-WordNet data shows a more pronounced enhancement compared to the random-WordNet data. The significant improvement in Area Match for the landmark-WordNet data is particularly noteworthy as area Match is a more granular metric that requires higher precision from the model. While unenhanced landmark data may sometimes perform less well than random data, the value of landmark data becomes evident when augmented with semantic resources like WordNet and the LLMs rewrite and RAG augmentation methods we later shown.
7.2 LLMs Rewrite and RAG augmentation
We then investigate the performance of large language models (LLMs) for rewriting and RAG-augmented data. This investigation aims to understand how these methods can enhance the quality and diversity of the training data, potentially leading to further improvements in model performance.
  • CoT+RAG: This method involves using a larger Qwen 2.5 72B model with RAG to provide more information and CoT prompt instructions to sequentially generate answers. The prompts used for this process are detailed in Table 3.3, and no parameter tuning was involved.
  • Rewrite Data: In this approach, the Landmark Data is rewritten using a Qwen 2.5 72B language model, as detailed in Table 3.2. This expands the dataset to 3,000 samples from Reuters and 6,000 from Web of Science (WoS).
  • RAG Augmented Data: In this approach, each document is augmented using a Qwen 2.5 72B language model, as detailed in Table 3.4. This expands the dataset to 15,473 samples from Reuters and 62,886 from WoS.
  • Combined Data: This method combines the WordNet, LLM rewriting, and RAG augmentation techniques by merging the training data together.
  • Original Data: For comparison only, we also list the performance of the model when using all training data with their true labels. The original datasets contain 5,394 true labels for Reuters and 23,492 true labels for WoS.
Dataset Part Match All Match In Right Order
Reuters, CoT+RAG 82.52% 73.03% 70.60%
Reuters, Landmark data 90.85% 82.35% 75.86%
Reuters, Landmark wordnet data 93.05% 83.97% 78.82%
Reuters, Rewrite Data 91.47% 83.73% 78.08%
Reuters, RAG Augment Data 95.14% 85.68% 78.87%
Reuters, Combined Augment Data 95.41% 89.02% 81.60%
Reuters, Original Data, for reference only 96.66% 91.89% 85.77%
Table 9: Detailed results for Reuters dataset
From the results presented in Table 9, it is clear that even with a larger model like Qwen 2.5 72B employed for CoT+RAG, its performance does not match that of a much smaller, parameter fine-tuned model such as Qwen 2.5 0.5B, which outperforms the larger model across multiple metrics. Specifically, for the Reuters dataset, the RAG Augmented Data method significantly outperformed other data augmentation techniques. Notably, despite using only 300 landmark labels, the RAG Augmented Data achieved a part match rate that was only 1.52% lower than the result obtained using the full set of 5,394 training data labels, which is a remarkable achievement. On the other hand, it is worth noting that when using only RAG augmented data, the accuracy for "All Match" is comparatively lower. This suggests that while RAG augmented data excel at partial matching, they may be more challenging to clearly differentiate for complete matches. Moreover, the Combined Augmentation method, which integrates WordNet, LLM rewriting, and RAG augmentation, achieved Part Match, All Match, and In Right Order scores of 95.41%, 89.02%, and 81.60%, respectively, on the Reuters dataset. This demonstrates that our strategy of integrating data with varying levels of vocabulary overlap has indeed contributed to improved metrics. The ensemble approach effectively leverages the strengths of different data augmentation techniques, leading to a more robust and accurate model, particularly for the "All Match" and "In Right Order" categories, thereby addressing the limitations observed with RAG data.
Domain Match Area Match
WOS, CoT+RAG 63.54% 43.69%
WOS, Landmark data 73.85% 47.48%
WOS, Landmark wordnet data 77.35 % 52.09%
WOS, Rewrite Data 75.46% 50.72%
WOS, RAG Augment Data 77.67% 52.67%
WOS, Combined Data 82.43% 60.02%
WOS, Original Data 90.32% 81.01%
Table 10: Detailed results for WOS dataset
From the results presented in Table 10, we observe a less pronounced performance improvement for RAG augmentation compared to the Reuters dataset. Specifically, even with the RAG Augmented Data method, the performance in terms of Domain Match and Area Match only marginally exceeds that of the Landmark WordNet data. This can be attributed to several factors: research topics are inherently more complex and specialized than news articles, making the LLM rewriting and RAG generation processes less effective compared to traditional WordNet methods, which typically involve replacing only certain "safe" words. Additionally, the terminology and concepts found in scientific literature are more technical and abstract, posing greater challenges for the model to accurately understand and generate high-quality text.
Despite these challenges, fine-tuned models consistently outperform the RAG-CoT approach without parameter tuning, highlights the importance of task-specific model fine-tuning. Moreover, it is clear that using all original training data, which includes over 20,000 labeled samples, leads to significantly better overall performance. However, in real-world scenarios where labeled data is often limited, our methods for generating additional augmented data still contribute to huge boost in performance improvements. This emphasizes the significance of data augmentation techniques in enhancing model performance, especially when sample sizes are constrained. Notably, combining the Landmark, rewrite, and RAG Augment data results in a substantial improvement, with an 8.58% increase in Domain Match and a 16.33% increase in Area Match compared to using only the landmark data. The significant improvement in area classification is particularly encouraging, as it suggests that the different levels of vocabulary overlap in the combined augmented data cover a broader range of case types, thereby leading to a significant performance boost when used together rather than individually. These findings not only emphasize the critical role of data augmentation in boosting model performance but also suggest that different strategies should be adopted for optimizing model performance across various types of text data.
8 Conclusion and Future Works
In this study, we have introduced an innovative framework for fine-tuning large language models (LLMs) in text classification tasks by integrating multiple advanced techniques: clustering for data selection, retrieval-augmented generation (RAG) for data augmentation, and chain-of-thought (CoT) reasoning. Our experimental results demonstrate that using clustering algorithms for data selection effectively identifies representative landmark data and finds near neighbors for RAG, significantly enhancing the model’s accuracy compared to random selection and other existing methods. This is particularly evident in datasets such as Reuters 20 Newsgroups and Web of Science, where the task involves classifying documents into over 100 categories. To the best of our knowledge, we are the first to combine clustering and RAG for text data augmentation. Our novel data augmentation approach, leveraging RAG and CoT, not only expands the dataset but also enhances its diversity and richness, leading to substantial improvements in classification accuracy. By integrating all stages of augmented data—from LLM rewrites to RAG and traditional similar word replacements with varying vocabulary overlap levels—our method ensures that the training data is both diverse and representative. This enables the model to handle various practical scenarios more proficiently, improving its generalization ability and accuracy. Furthermore, our methods are straightforward to implement and rely on open-source tools, making them highly practical and accessible. Overall, our study provides a robust and efficient solution for text classification tasks, offering significant potential for widespread adoption and application. For future research, several areas can be explored to further improve our framework. One direction is to implement constrained decoding to ensure that LLMs output only the available categories and adhere to specific regular expressions. This can be crucial for achieving 100% usability in practical applications. Another area is to explore different methods for data augmentation and compare them with our current approach. While we used LLMs to select landmarks, other methods based on numerical metrics might prove more effective. Additionally, scaling up the model to larger LLMs and comparing their performance with even larger models, such as QLORA 70B model, will provide valuable insights into the scalability and efficiency of our framework.